v=4/3pi^2h for h
what is the solution to this problem??

Answer by venugopalramana(3286)

ANSWER TO YOU PROBLEM IS GIVEN BELOW
-----------------------------------------------------------------------------
V=4(PIE)^2*H/3...MULTIPLY BOTH SIDES WITH 3
V*3=3*4(PIE)^2*H/3=4(PIE)^2*H.....NOW DIVIDE OTH SIDES WITH 4(PIE)^2
3V/(4(PIE)^2)=4(PIE)^2*H/(4(PIE)^2)=H
HENCE H=3V/(4(PIE)^2)
-----------------------------------------------------------------------------------------------------------